# SVIP: Semantically Contextualized Visual Patches for Zero-Shot Learning

### v1

@@ Abstract
Zero-shot learning (ZSL) aims to recognize unseen classes without labeled training examples by leveraging class-level semantic descriptors such as attributes. A fundamental challenge in ZSL is semantic misalignment, where semantic-unrelated information involved in visual features introduce ambiguity to visual-semantic interaction. Unlike existing methods that suppress semantic-unrelated information post hoc either in the feature space or the model space, we propose addressing this issue at the input stage, preventing semantic-unrelated patches from propagating through the network. To this end, we introduce Semantically contextualized VIsual Patches (SVIP) for ZSL, a transformer-based framework designed to enhance visual-semantic alignment. Specifically, we propose a self-supervised patch selection mechanism that preemptively learns to identify semantic-unrelated patches in the input space. This is trained with the supervision from aggregated attention scores across all transformer layers, which estimate each patch’s semantic score. As removing semantic-unrelated patches from the input sequence may disrupt object structure, we replace them with learnable patch embeddings. With initialization from word embeddings, we can ensure they remain semantically meaningful throughout feature extraction. Extensive experiments on ZSL benchmarks demonstrate that SVIP achieves state-of-the-art performance results while providing more interpretable and semantically rich feature representations.
@@ 1 Introduction
Specifically, to mitigate the semantic misalignment problem comprehensively, we propose self-supervised patch selection mechanism that learns to identify the semantic-unrelated patches in the input space, i.e., before feeding into the transformer blocks. In particular, we aggregate the self-attention weight matrices across all transformer blocks, yielding reliable semantic understanding about the patch importance throughout the entire ViT. To mitigate the negative impact of semantic-unrelated patches to semantic-related patches, we train a patch classifier to determine the input patches that are semantic-unrelated patches, using the supervision from the aggregated attention matrix. At the test time, the patch classifier can directly prevent the semantic-unrelated patches from propagating into the feature extraction. Once the semantic-unrelated patches are identified in the input sequence, a naive solution is removing them from the input sequence. However, this can break object structure on the image. Instead, we propose to mask the semantic-unrelated patches with a learnable semantic context patch. The semantic context patch is initialized with the attribute-level word embeddings that provides rich semantic information. These contextualize input patches can enhance the visual-semantic interaction throughout the visual feature extraction process.
Overall, the contributions of this paper are:
@@
  • We propose a self-supervised patch selection strategy to identify semantic-unrelated patches using the attention scores aggregated from the self-attention weight matrix across the ViT blocks. We propose patch semantic contextualization that injects of attribute-level word embeddings into semantic-unrelated patches, ensuring that semantic information is incorporated into visual representations from the initial layers of the transformer.
@@ 2 Related Work
The core objective in ZSL [38, 47, 23, 49, 8, 6, 41, 7, 13, 20, 19, 10] is to generalize visual-semantic knowledge obtained from seen classes to unseen classes. This generalization relies on semantic descriptors, which are typically represented as attributes [18], documents [17], or word embeddings [32]. The main stream ZSL methods can be generally categorized as generative methods and embedding-based methods. Generative ZSL methods typically synthesize unseen visual features from semantic descriptors using generative models, enabling standard classification pipelines to incorporate synthetic training samples [49, 51, 8]. These approaches leverage various generative architectures, including GANs [49, 54, 41], VAEs [46, 39] and Normalizing Flows [40, 11, 9]. Embedding-based ZSL methods aim to learn a direct mapping from the visual to the semantic space, aligning image representations with class descriptors. These methods typically leverage the attention mechanism to localize the semantic descriptors. Zhu et al. [55] introduced a semantic-guided multi-attention localization network to extract the local fine-grained features of the discriminative patches. Xu et al. [50] proposed an attribute prototype network, where each prototype represents a specific semantic concept and is used as the 1×1 filter to convolve the feature maps with the feature maps from ResNet101 in order to attend to the region of interest. The follow-up work [28] proposed to use the GloVe [35] model to extract semantic vectors from the attribute names (e.g., “white breast”), which are further used to represent the attribute prototypes. More recently, Chen et al. [3] developed a mutually semantic distillation network to progressively refine the alignment between visual and semantic features.
@@ 3.5 Attribute Localization
Although the final class token ${\bm{v}}_{cls}$ can capture the global representation, the semantic-related patches can help localize attributes. Let ${\bm{Z}}^{L^{\prime}}=[{\bm{z}}_{i}^{{}^{\prime}}|i\in\mathcal{S}_{M}]$ be the final representations of the top $M$ semantic-related patches. We learn a Patch-to-Attribute projection function (P2A) to map the patch embeddings ${\bm{Z}}^{L^{\prime}}\in\mathbb{R}^{M\times C}$ into the attribute space: $\hat{{\bm{A}}}=\texttt{P2A}({\bm{Z}}^{L^{\prime}})\in\mathbb{R}^{M\times K}$, where $K$ is the number of attributes the describes a category. A max pooling operation is then applied over these patch-attribute embeddings:
@@ 3.6 Model Optimization and Inference
ResNet101 TF [31] ECCV’20 64.9 52.8 64.7 58.1 72.2 59.8 75.1 66.6 66.0 45.6 40.7 43.0
RFF [22] CVPR’20 - 52.6 56.6 54.6 - 59.8 75.1 66.5 - 45.7 38.6 41.9
APN [50] NeurIPS’20 72.0 65.3 69.3 67.2 68.4 56.5 78.0 65.5 61.6 41.9 34.0 37.6
@@ 4.2 Main Results
Comparison with CNN Backbones. We next first compare SVIP with a group of methods that adopt ResNet101 as their backbone (upper block of Table 1). Specifically, approaches such as APN [50], GEM [28], MSDN [3], Dis‐VAE, and ICCE all operate on global CNN‐extracted features to align visual and semantic spaces for zero‐shot learning. Despite their efficacy on certain datasets, our SVIP achieves consistently state-of-the-art performance results across CUB, AwA2, and SUN. Notably, SVIP obtains an H value of 76.1% on CUB, 75.5% on AwA2, and 51.2% on SUN, surpassing all competing ResNet101‐based methods. These results demonstrate SVIP’s effectiveness in capturing fine‐grained semantic cues while maintaining strong discriminative ability for both seen (S) and unseen (U) classes.
@@
Comparison with ViT Backbones. We next compare SVIP with recent approaches with ViT as backbone, including language-driven methods (CLIP [37], CoOp [53], I2DFormer [29], I2MV [30]), attribute-based methods (DUET[12], ZSLViT [4]). Generally, ViT‐based methods benefit from attention mechanisms over patch embeddings, but often lack explicit supervision for semantic‐unrelated patches. As shown, SVIP delivers the best performance in terms of unseen accuracy (U) on CUB (72.8%) and SUN (54.2%), as well as a leading harmonic mean on both datasets. On AwA2, our approach also shows competitive results, attesting to the adaptability of SVIP’s localized attribute reasoning across a range of visual domains.
@@
Ablation Study. We conduct an ablation study to analyze the contribution of each component in SVIP (Table2). The Baseline (ViT) uses a simple linear layer on the class token without patch-level operations, yielding the lowest performance. SVIP w/o SSPS removes the Self-Supervised Patch Selection, applying semantic patches indiscriminately, leading to a significant drop, highlighting the necessity of targeted patch selection. SVIP w/o PSC directly removes semantic-unrelated patches instead of contextualizing them, which slightly disrupts object structure and lowers performance. SVIP w/o JSD excludes Jensen–Shannon divergence, reducing training stability and overall accuracy. SVIP w/o W2P replaces word embeddings with randomly initialized patch embeddings, resulting in degraded semantic alignment. SVIP w/o P2A removes Patch-to-Attribute mapping, forcing reliance on the class token, which weakens unseen class accuracy. The full SVIP model achieves the best results across all datasets, demonstrating that each module contributes to improving visual-semantic alignment and overall ZSL performance.
@@
Hyper-parameter sensitivity. In a series of experiments, we investigate the influence of the four hyper-parameters on the performance of our proposed method. Figure 4(a) shows the effect of varying the number of preserved patches $M$ from 49 to 35. As the dropped patch number increases, the overall performance also improves. However, the performance saturation occurs at around 40, and larger patch number tend to decrease performance due to over-pruned semantic-related patches. Figure 4(b) demonstrates that the JS Divergence is a considerably sensitive hyper-parameter. We observe that set this number to 1 achieves the best performance. In Figure 4(c), we vary the loss coefficient for patch classfier $\lambda_{2}$ and found that both ZSL and GZSL achieve the best performance at 3. In Figure 4(d), we found that the temperature scale has a considerable impact on the performance variation, which peaks at 5.
@@
Figure 6: (a) Example images with semantic‐unrelated patches masked. (b) t-SNE visualization of the image patches, where red bounding boxes indicate the semantical-unrelated patches.
@@ 4.3 Qualitative Study
Attribute Localization. We present attention maps (i.e., the layer before max pooling operation) in Figure 5, for randomly selected attributes on three example images from the CUB dataset. The attention maps are generated by applying max pooling operations to the model output, highlighting the regions where the model is focusing on each attribute. The regions that are masked with striped hatch are identified as semantic-unrelated patches and thus not considered for attribute localization. The baseline method ViT focuses on many incorrect patches where are classified as semantic-unrelated patches in SVIP, which demonstrates the usefulness of SSPS.

### v2

@@ Abstract
Zero-shot learning (ZSL) aims to recognize unseen classes without labeled training examples by leveraging class-level semantic descriptors such as attributes. A fundamental challenge in ZSL is semantic misalignment, where semantic-unrelated information involved in visual features introduce ambiguity to visual-semantic interaction. Unlike existing methods that suppress semantic-unrelated information post hoc either in the feature space or the model space, we propose addressing this issue at the input stage, preventing semantic-unrelated patches from propagating through the network. To this end, we introduce Semantically contextualized VIsual Patches (SVIP) for ZSL, a transformer-based framework designed to enhance visual-semantic alignment. Specifically, we propose a self-supervised patch selection mechanism that preemptively learns to identify semantic-unrelated patches in the input space. This is trained with the supervision from aggregated attention scores across all transformer layers, which estimate each patch’s semantic score. As removing semantic-unrelated patches from the input sequence may disrupt object structure, we replace them with learnable patch embeddings. With initialization from word embeddings, we can ensure they remain semantically meaningful throughout feature extraction. Extensive experiments on ZSL benchmarks demonstrate that SVIP achieves state-of-the-art performance results while providing more interpretable and semantically rich feature representations. Code is available at https://github.com/uqzhichen/SVIP.
@@ 1 Introduction
Specifically, to mitigate the semantic misalignment problem comprehensively, we propose self-supervised patch selection mechanism that learns to identify the semantic-unrelated patches in the input space, i.e., before feeding into the transformer blocks. In particular, we aggregate the self-attention weight matrices across all transformer blocks, yielding reliable semantic understanding about the patch importance throughout the entire ViT. To mitigate the negative impact of semantic-unrelated patches to semantic-related patches, we train a patch classifier to determine the input patches that are semantic-unrelated patches, using the supervision from the aggregated attention matrix. At the test time, the patch classifier can directly prevent the semantic-unrelated patches from propagating into the feature extraction. Once the semantic-unrelated patches are identified in the input sequence, a naive solution is removing them from the input sequence. However, this can break object structure on the image. Instead, we propose to mask the semantic-unrelated patches with a learnable semantic context patch. The semantic context patch is initialized with the attribute-level word embeddings that provides rich semantic information. These contextualize input patches can enhance the visual-semantic interaction throughout the visual feature extraction process. Overall, the contributions of this paper are:
@@
  • We propose a self-supervised patch selection strategy to identify semantic-unrelated patches using the attention scores aggregated from the self-attention weight matrix across the ViT blocks. We propose patch semantic contextualization that injects attribute-level word embeddings into semantic-unrelated patches, ensuring that semantic information is incorporated into visual representations from the initial layers of the transformer.
@@ 2 Related Work
In the deep learning era [62, 5, 53, 63, 57, 56, 50, 31, 69, 16, 15, 65, 70, 68, 71, 67, 54, 52, 55], the core objective in ZSL [42, 58, 28, 60, 9, 7, 45, 8, 14, 25, 24, 11, 22, 23] is to generalize visual-semantic knowledge obtained from seen classes to unseen classes. This generalization relies on semantic descriptors, which are typically represented as attributes [21], documents [20], or word embeddings [36]. The mainstream ZSL methods can be generally categorized as generative methods and embedding-based methods. Generative ZSL methods typically synthesize unseen visual features from semantic descriptors using generative models, enabling standard classification pipelines to incorporate synthetic training samples [60, 64, 9]. These approaches leverage various generative architectures, including GANs [60, 73, 45], VAEs [51, 43] and Normalizing Flows [44, 12, 10]. Embedding-based ZSL methods aim to learn a direct mapping from the visual to the semantic space, aligning image representations with class descriptors. These methods typically leverage the attention mechanism to localize the semantic descriptors. Zhu et al. [74] introduced a semantic-guided multi-attention localization network to extract the local fine-grained features of the discriminative patches. Xu et al. [61] proposed an attribute prototype network, where each prototype represents a specific semantic concept and is used as the 1×1 filter to convolve the feature maps with the feature maps from ResNet101 in order to attend to the region of interest. The follow-up work [33] proposed to use the GloVe [39] model to extract semantic vectors from the attribute names (e.g., “white breast”), which are further used to represent the attribute prototypes. More recently, Chen et al. [3] developed a mutually semantic distillation network to refine the alignment between visual and semantic features.
@@ 3.5 Attribute Localization
Although the final class token ${\bm{v}}_{cls}$ can capture the global representation, the semantic-related patches can help localize attributes. Let ${\bm{Z}}^{L^{\prime}}=[{\bm{z}}_{i}^{{}^{\prime}}|i\in\mathcal{S}_{M}]$ be the final representations of the top $M$ semantic-related patches. We learn a Patch-to-Attribute projection function (P2A) to map the patch embeddings ${\bm{Z}}^{L^{\prime}}\in\mathbb{R}^{M\times C}$ into the attribute space: $\hat{{\bm{A}}}=\texttt{P2A}({\bm{Z}}^{L^{\prime}})\in\mathbb{R}^{M\times K}$, where $K$ is the number of attributes that describes a category. A max pooling operation is then applied over these patch-attribute embeddings:
@@ 3.6 Model Optimization and Inference
ResNet101 RFF [27] CVPR’20 - 52.6 56.6 54.6 - 59.8 75.1 66.5 - 45.7 38.6 41.9
CANZSL [6] WACV’20 60.6 47.9 58.1 52.5 68.9 49.7 70.2 58.2 60.1 46.8 35.0 40.0
@@ 4.2 Main Results
Comparison with CNN Backbones. We next first compare SVIP with a group of methods that adopt ResNet101 as their backbone (upper block of Table 1). Specifically, approaches such as APN [61], GEM [33], MSDN [3], Dis‐VAE, and ICCE all operate on global CNN‐extracted features to align visual and semantic spaces for zero‐shot learning. Despite their efficacy on certain datasets, our SVIP achieves consistently state-of-the-art performance results across CUB, AwA2, and SUN. Notably, SVIP obtains an H value of 75.0% on CUB, 74.9% on AwA2, and 50.7% on SUN, surpassing all competing ResNet101‐based methods. These results demonstrate SVIP’s effectiveness in capturing fine‐grained semantic cues while maintaining strong discriminative ability for both seen (S) and unseen (U) classes.
@@
Comparison with ViT Backbones. We next compare SVIP with recent approaches with ViT as backbone, including language-driven methods (CLIP [41], CoOp [72], I2DFormer [34], I2MV [35]), attribute-based methods (DUET[13], ZSLViT [4]). Generally, ViT‐based methods benefit from attention mechanisms over patch embeddings, but often lack explicit supervision for semantic‐unrelated patches. As shown, SVIP delivers the best performance in terms of unseen accuracy (U) on CUB (72.1%) and SUN (53.7%), as well as a leading harmonic mean on both datasets. On AwA2, our approach also shows competitive results, attesting to the adaptability of SVIP’s localized attribute reasoning across a range of visual domains.
@@
Ablation Study. We conduct an ablation study to analyze the contribution of each component in SVIP (Table 2). The Baseline (ViT) uses a simple linear layer on the class token without patch-level operations, yielding the lowest performance. SVIP w/o SSPS removes the Self-Supervised Patch Selection, applying semantic patches indiscriminately, leading to a significant drop, highlighting the necessity of targeted patch selection. SVIP w/o PSC directly removes semantic-unrelated patches instead of contextualizing them, which slightly disrupts object structure and lowers performance. SVIP w/o JSD excludes Jensen–Shannon divergence, reducing training stability and overall accuracy. SVIP w/o W2P replaces word embeddings with randomly initialized patch embeddings, resulting in degraded semantic alignment. SVIP w/o P2A removes Patch-to-Attribute mapping, forcing reliance on the class token, which weakens unseen class accuracy.
@@
Hyper-parameter sensitivity. In a series of experiments, we investigate the influence of the four hyper-parameters on the performance of our proposed method. Figure 4 (a) shows the effect of varying the number of preserved patches $M$ from 49 to 35. The performance saturation occurs at around 40, and larger patch numbers tend to decrease performance due to over-pruned semantic-related patches. Figure 4 (b) demonstrates that the JS Divergence is a considerably sensitive hyperparameter, achieving the best performance at 1. In Figure 4 (c), we vary the loss coefficient for patch classfier $\lambda_{2}$ and found that both ZSL and GZSL achieve the best performance at 3. In Figure 4 (d), we found that the temperature scale has a considerable impact on the performance variation, which peaks at 5.
@@
Figure 6: (a) Example images with semantic‐unrelated patches masked. (b) t-SNE visualization of the image patches, where red bounding boxes indicate the semantically unrelated patches.
@@ 4.3 Qualitative Study
Attribute Localization. We present attention maps (i.e., the layer before max pooling operation) in Figure 5, for randomly selected attributes on three example images from the CUB dataset. The attention maps are generated by applying max pooling operations to the model output, highlighting the regions where the model is focusing on each attribute. The regions that are masked with striped hatch are identified as semantic-unrelated patches and thus not considered for attribute localization. The baseline method ViT focuses on many incorrect patches that are classified as semantic-unrelated patches in SVIP, which demonstrates the usefulness of SSPS.
@@ 5 Conclusions
Acknowledgment
This work is supported by the Australian Research Council under the streams of Discovery Project (No. DP240101814).